OpenStudy (courtney_celiene1236):

prove using mathematical induction 12 + 42 + 72 + ... + (3n - 2)2 = (n(〖6n〗^2-3n-1))/2

OpenStudy (alivejeremy):

What?

OpenStudy (p0sitr0n):

start with the first term, then your IH will be that if that holds for n, it must hold for n+1 also (f(n)=g(n) => f(n+1)=g(n+1))

OpenStudy (courtney_celiene1236):

Here's what I have so far: \[s _{1}=\frac{ n(6n ^{2}-3n-1) }{2}\] \[=\frac{ 1(6(1^{2})^{2}-3(1)-1) }{ 2 }\] \[=\frac{ 1(2) }{ 2 }\] \[S _{1}=1\]

OpenStudy (courtney_celiene1236):

@angela101

OpenStudy (mathmate):

@courtney_celiene1236 Good, that's a good first step. By the way, we at OpenStudy here write \(1^2\) as 1^2. It's ugly, but more understandable than 12. If we denote: \(f(x)=1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 \) and \(\large g(x)=\frac{n(6n^2-3n-1)}{2}\) then the next step you need to do is to show that "given f(n)=g(n), then g(n+1)=f(n+1)" ...............(1) this is the same as proving g(n+1)-g(n)=f(n+1)-f(n) .....................(2) From the definition of f(n), we know that \(f(n+1)=f(n)+(3(n+1)-2)^2\) which means that \(f(n+1)-f(n)=(3(n+1)-2)^2\) that reduces (2) to proving \(g(n+1)-g(n)=(3(n+1)-2)^2\) ................(3) So expanding (3) is all you need to complete the proof.